# Mixture of Experts with Mixture of Precisions for Tuning Quality of Service

### v1

@@ Abstract
The increasing demand for deploying large Mixture-of-Experts (MoE) models in resource-constrained environments necessitates efficient approaches to address their high memory and computational requirements challenges. Moreover, given that tasks come in different user-defined constraints and the available resources change over time in multi-tenant environments, it is necessary to design an approach which provides a flexible configuration space. This paper presents an adaptive serving approach for the efficient deployment of MoE models, capitalizing on partial quantization of the experts. By dynamically determining the number of quantized experts and their distribution across CPU and GPU, our approach explores the Pareto frontier and offers a fine-grained range of configurations for tuning throughput and model quality. Our evaluation on an NVIDIA A100 GPU using a Mixtral 8x7B MoE model for three language modelling benchmarks demonstrates that the throughput of token generation can be adjusted from 0.63 to 13.00 token per second. This enhancement comes with a marginal perplexity increase of 2.62 to 2.80, 6.48 to 7.24, and 3.24 to 3.53 for WikiText2, PTB, and C4 datasets respectively under maximum quantization. These results highlight the practical applicability of our approach in dynamic and accuracy-sensitive applications where both memory usage and output quality are important.
@@ 4.2 Results
Here, we first evaluate the effect of quantized experts on the performance of the MoE model. As shown in Figure 2, increasing the number of 4-bit experts generally leads to a decrease in the quality of the generated output (higher perplexity), although this trend is not strictly monotonic. In the case of the WikiText2 and C4 datasets, there are points in the graph where increasing the number of quantized experts decreases the perplexity. This might be due to the perplexity metric’s limitation in accurately evaluating minor modifications to the MoE model and also quantization smoothing the decision boundaries of the model [32].
@@ 5 Conclusion
In this paper, we propose an adaptive serving approach for efficient deployment of large MoE models in dynamic single-GPU constrained settings. Our approach enhances performance by reducing the overall memory footprint of MoE models through partial expert quantization. By employing partial expert quantization, our approach offers a fine-grained configuration space, allowing precise control over both throughput and output quality. Therefore, as user-defined constraints and available resources change over time, our adaptive serving approach is able to reconfigure the deployment plan and adjust to the new settings. We evaluate our adaptive serving approach on an NVIDIA A100 GPU using a Mixtral 8x7B MoE model for three language modeling benchmarks. Our results show that under GPU memory constraints of 26.28 to 53.03 GB, a throughput of 0.63 to 13.00 tokens per second is achievable. We also demonstrate that, in terms of output quality, keeping the non-expert layers in 16-bit and partially quantizing the experts to 4-bit results in perplexity ranges of 2.55 to 2.80, 6.47 to 7.24, and 3.21 to 3.53 for the WikiText2, PTB, and C4 datasets, respectively.

### v2

@@ Abstract
The increasing demand for deploying large Mixture-of-Experts (MoE) models in resource-constrained environments necessitates efficient approaches to address their high memory and computational requirements challenges. Moreover, given that tasks come in different user-defined constraints and the available resources change over time in multi-tenant environments, it is necessary to design an approach which provides a flexible configuration space. This paper presents an adaptive serving approach for the efficient deployment of MoE models, capitalizing on partial quantization of the experts. By dynamically determining the number of quantized experts and their distribution across CPU and GPU, our approach explores the Pareto frontier and offers a fine-grained range of configurations for tuning throughput and model quality. Our evaluation on an NVIDIA A100 GPU using a Mixtral 8x7B MoE model for three language modelling benchmarks demonstrates that the throughput of token generation can be adjusted from 0.63 to 13.00 token per second. This enhancement comes with a marginal perplexity increase of 3.81 to 4.00, 13.59 to 14.17, and 7.24 to 7.40 for WikiText2, PTB, and C4 datasets respectively under maximum quantization. These results highlight the practical applicability of our approach in dynamic and accuracy-sensitive applications where both memory usage and output quality are important.
@@ 4.2 Results
Here, we first evaluate the effect of quantized experts on the performance of the MoE model. As shown in Figure 2, increasing the number of 4-bit experts generally leads to a decrease in the quality of the generated output (higher perplexity), although this trend is not always strictly monotonic. In the case of the PTB dataset, there are points in the graph where increasing the number of quantized experts decreases the perplexity. This might be due to the perplexity metric’s limitation in accurately evaluating minor modifications to the MoE model and also quantization smoothing the decision boundaries of the model [32].
@@ 5 Conclusion
In this paper, we propose an adaptive serving approach for efficient deployment of large MoE models in dynamic single-GPU constrained settings. Our approach enhances performance by reducing the overall memory footprint of MoE models through partial expert quantization. By employing partial expert quantization, our approach offers a fine-grained configuration space, allowing precise control over both throughput and output quality. Therefore, as user-defined constraints and available resources change over time, our adaptive serving approach is able to reconfigure the deployment plan and adjust to the new settings. We evaluate our adaptive serving approach on an NVIDIA A100 GPU using a Mixtral 8x7B MoE model for three language modeling benchmarks. Our results show that under GPU memory constraints of 26.28 to 53.03 GB, a throughput of 0.63 to 13.00 tokens per second is achievable. We also demonstrate that, in terms of output quality, keeping the non-expert layers in 16-bit and partially quantizing the experts to 4-bit results in perplexity ranges of 3.81 to 4.00, 13.59 to 14.17, and 7.24 to 7.40 for the WikiText2, PTB, and C4 datasets, respectively.